two particles a and b are released from rest, having equal charges of -8 x10-6 c and initially separated from each other by 0.025m. the masses of particles a and b are 6.25 x10-6 kg and 3.75x10-6 kg respectively. find the magnitude of the initial accelerations of particles a and b.


1
Expert's answer
2022-02-20T08:20:10-0500

Given:

"q_1=q_2=\rm -8 *10^{-6}\: C"

"m_1=\rm 6.25*10^{-6}\: kg"

"m_2=\rm 3.75*10^{-6}\: kg"

"r=0.025\:\rm m"


The Coulomb's law says



"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9\frac{(8.0*10^{-6})^2}{0.025^2}=921.6 \:\rm N"


The acceleration of particles


"a_1=\frac{F}{m_1}=\frac{921.6}{6.25*10^{-6}}=1.47*10^8\:\rm m/s^2""a_2=\frac{F}{m_2}=\frac{921.6}{3.75*10^{-6}}=2.46*10^8\:\rm m/s^2"